OpenStudy (anonymous):

How long will it take a man to drive from Town A to town B 120 away and back if he drives to Town B at a speed of 40km/h and back to Town A at a speed of 8km/h faster?

OpenStudy (anonymous):

so 120 km there and 120 km back so time to get there: v=40km/h v=d/t; t=d/v t=120km/40km/h t=3 hours time to get back: v=48km/h t=120km/48km/h t=2.5 hours total time = time there plus time back total time = 3 hours +2.5 hours = 5.5 hours t=5.5 hours .....................final ans.